Using Bohr’s theory, calculate the energy required to remove the electron from the ground state of singly ionized helium, that is, helium with one electron already removed.


1
Expert's answer
2021-12-20T10:29:56-0500

The energy is


"E=\frac{13.6Z^2}{n^2}=\frac{13.6·2^2}{1^2}=54.4\text{ eV}."